### v1

# T-MAE: Temporal Masked Autoencoders for Point Cloud Representation Learning

Abstract
The scarcity of annotated data in outdoor point cloud segmentation poses a significant obstacle in harnessing the modeling capabilities of advanced networks like transformers. Consequently, scholars have been actively investigating efficacious self-supervised pre-training strategies, e.g. contrasting learning and reconstruction-based pretext tasks. Nevertheless, temporal information, which is inherent in the LiDAR point cloud sequence, is consistently disregarded. To better utilize this property, we propose an effective pre-training strategy, namely Temporal Masked Auto-Encoders (T-MAE), which takes as input temporally adjacent frames and learns temporal dependency. A SiamWCA backbone, containing a Siamese encoder and a window-based cross-attention (WCA) module, is established for the two-frame input. Taking into account that the motion of an ego-vehicle alters the illumination angles of the same instance, temporal modeling also serves as a robust and natural data augmentation, enhancing the comprehension of target objects. Moreover, instead of utilizing consecutive frames, it is more cost-effective and powerful by using distant historical frames. SiamWCA is a powerful architecture but heavily relies on annotated data. With our T-MAE pre-training strategy, we achieve the best performance on the Waymo dataset among self-supervised learning methods. Comprehensive experiments are conducted to validate all components of our proposal. Upon acceptance, the source code will be made accessible.
1 Introduction
Figure 1: T-MAE performance on Waymo dataset [39]. Our T-MAE pre-training achieves substantially better results than random initialization. It also achieve better results than the SOTA SSL method MV-JAR [48] with significantly fewer iterations. Each point triplet corresponds to a specific fine-tuned model variant and shows the performance difference with two arrows. From left to right, the triplets correspond to a model fine-tuned with 5%, 10% and 20% data.
In the domains of language and vision, researchers have observed that as models become more complex, the available amount of labeled data is often insufficient to adequately train huge models [11, 18] e.g. Vision Transformer (ViT) [12]. Consequently, there is a growing interest in exploring self-supervised learning (SSL) approaches as a potential solution to overcome this limitation. SSL serves as a pre-training technique with unlabeled data, accelerating the convergence of the models and improving their performance for downstream tasks [16, 6, 9]. The same challenge extends to the domain of point clouds, where annotations are more costly and time-consuming to obtain [46, 14]. For instance, a mere 10% and 0.8% of sweeps are annotated in the nuScene [2] and ONCE [29] datasets. This challenge makes pre-training for point cloud understanding a non-trivial endeavor.
Prior works mainly focus on synthetic and isolated objects [50, 56, 54, 35, 27] and indoor scene understanding [19, 24, 49, 36]. Transferring these methods to outdoor LiDAR points is challenging due to their sparsity and dynamic environmental conditions. At present, most self-supervised methods used for understanding point clouds in autonomous driving rely on contrastive learning [47, 53, 32, 34, 33]. These approaches model the similarity and dissimilarity between entities, such as segments [33, 45] and/or points [47]. In the wake of masked image modeling as a pretext task [18], efforts have also been devoted to the reconstruction of masked points [31, 41, 52, 48]. The main idea is randomly masking points or voxels and urging the network to infer the coordinates of points [52] and/or voxels [48] or other properties, e.g. occupancy [31, 1] and curvature [41]. Nevertheless, these methodologies often operate within the confines of a single-frame scenario, disregarding the fact that LiDAR data is typically acquired on a frame-by-frame basis. In other words, the valuable semantic information that is introduced by temporally adjacent frames is often not exploited.
Several methods attempt to leverage temporal information [21, 45, 33, 25] by incorporating multi-frame input during the self-supervised phase but their core concepts remain grounded in contrastive learning. Specifically, the scans captured at different times are treated as augmented samples of the same scene, without including temporal correspondence into the modeling procedure.
Therefore, we propose a new self-supervised method, namely T-MAE, to exploit the accumulated observations. The key idea of T-MAE is to learn a powerful representation for point cloud scans and integrate features from the past for present use. During the pre-training stage, the current scan is voxelized with a high masking ratio, while the previous scan is fed entirely to the encoder. Then, the reconstruction pretext task is performed by incorporating three types of information: voxel embeddings of the past scan, visible voxel embeddings of the current scan, and the position of masked voxels. In this way, the WCA module learns to incorporate historical information into the current frame using unlabeled data. The proposed T-MAE pre-training strategy endows the network with both a powerful representation for sparse point clouds and the capacity to strengthen the present by learning from the past. As shown in Fig. 1, the proposed T-MAE achieves higher overall and pedestrian-specific mAPH than the randomly initialized baseline, namely the same model but trained from scratch. Moreover, T-MAE also outperforms state-of-the-art MV-JAR [48] with over 1.6×1.6$\times$1.6 × fewer training iterations. Our contributions are summarized as follows:
  • We propose T-MAE, a novel and effective SSL approach for sparse point clouds, that learns temporal modeling in the process of reconstructing masked points.
  • We design a SiamWCA backbone, containing a Siamese encoder and a window-based sparse cross-attention (WCA) module, to incorporate historical information.
  • Rather than concatenating multiple consecutive point cloud scans, our approach demonstrates that incorporating more distant historical frames provides greater advantages.
  • Our experiments demonstrate the efficacy of T-MAE by attaining impressive results on the Waymo Open dataset. T-MAE achieves 55.71 mAPH for pedestrians with only 5% of the labeled data and surpasses the SOTA SSL approaches by a substantial margin across all levels of fine-tuning data amount.
2 Related Work
2.1 Static Self-supervised Learning
SSL for point clouds is a burgeoning field due to the scarcity of annotations. In general, the pipeline of SSL consists of two phases. During the initial stage, namely pre-training with unlabeled data, the network is trained using a pretext task. Subsequently, in downstream tasks such as segmentation and detection, the pre-trained weights are loaded to the backbone, and the backbone is attached with task-specific heads. The resulting model is further fine-tuned using annotated data.
The initial research efforts are primarily directed towards the exploration of contrastive learning. The underlying hypothesis is that an image or 3D scene demonstrates feature equivalence even after undergoing different transformations [9, 5]. One of the key challenges in the point cloud domain is establishing correspondences across scenes. PointContrast [47] and DepthContrast [57] track the points while performing different transformations. GCC-3D [25] exploit sequential information to obtain pseudo instances and then perform contrasting on the instance level. SegContrast [32] firstly obtains segments by an unsupervised clustering and then contrasting segments between two transformed views. These methods allow the network to learn equivalence with regard to geometric transformations. However, contrastive learning-based SSL approaches usually suffer from careful tuning of hyperparameters and complicated pre- or post-processing to find correspondences.
The focus has shifted to reconstructing masked points, following the success of masked image modeling [18, 44]. The masking strategies remain consistent, i.e. voxelize a point cloud into cubes followed by random masking of these cubes. The reconstruction targets vary across papers. OccupancyMAE [31] classifies voxels to whether they are occupied. Geo-MAE [41] infers occupancy as well as the normal and the curvature of each masked voxel. The GD-MAE’s pretext task is to reconstruct a fixed number of points for masked voxels [52]. MV-JAR [48] partitions the masked Voxel into two categories, necessitating the network to either predict the coordinates of the voxels or produce the points themselves. In this paper, we employ the same masking strategy as GD-MAE [52] but expand the knowledge source to include both the current visible voxels and voxels from a previous frame as a reference.
2.2 Temporal Self-supervised Learning
There are many SSL methods designed to address temporal or spatiotemporal tasks. In video-related tasks, such as MAE-ST [15], a random patch masking strategy is employed over consecutive frames. The model is tasked with recovering these masked patches while considering information from neighboring frames. VideoMAE [42] retains the same reconstruction pipeline as the pretext task but exploits a tube masking strategy. SiameseMAE [17] learns object-centered representations with the help of cross-attention layers and an asymmetrical masking technique on consecutive frames. There are also content-based masking strategies, e.g. motion-guided masking [20, 30]. The primary concept underlying these methods is acquiring an understanding of temporal dependency across frames through the process of reconstructing patches with reference to consecutive frames.
While this concept is effectively used in video understanding, it is rarely applied to the learning of sparse point cloud representations. This is because point clouds are typically handled on a frame-by-frame basis rather than being regarded as a temporal sequence. For instance, when using temporally adjacent scans as input during self-supervised learning, both STSSL [45] and TARL [33] use HDBSCAN [3] to obtain segments and apply self-supervision by minimizing the feature distance between the same segments or points between neighbor frames. While they achieve impressive results, two issues persist 1) The use of elaborate and time-consuming pre-processing methods to obtain segments. 2) A primary focus on object consistency across frames rather than understanding object motion.
Unlike other approaches, this paper aims to reduce the reliance on complex preprocessing techniques and emphasizes the significance of enabling the model to establish temporal correspondence through self-supervised learning from unlabeled data.
3 Method
Figure 2: Overview of the proposed T-MAE pre-training. Two frames are sampled from a sequence of point clouds according to temporal batch-based strategy. The prior scan $\mathcal{P}^{t+t_{1}}$ is voxelized and directly fed into the Siamese windowed cross-attention (SiamWCA) backbone. The current scan $\mathcal{P}^{t+t_{2}}$ undergoes a masking process before fed into the backbone. Furthermore, pillar-wise tokens are computed by a Siamese hierarchical encoder. The two-way gray arrow indicates weight sharing. The WCA module takes as input the full tokens of the previous frame and the partial observation of the current frame and creates enhanced tokens. In the diffusion process, masked tokens are updated by their neighbor tokens from WCA by a group of convolution layers. Subsequently, a reconstruction head recovers masked points. During fine-tuning, the full SiamWCA backbone is initialized with pre-trained weights obtained from the pre-training phase.
In this section, previous works are briefly discussed in Sec. 3.1. Then, the proposed T-MAE pre-training strategy is demonstrated in Sec. 3.2. Eventually, the window-based sparse cross-attention (WCA) module is elaborated in Sec. 3.3.
3.1 Preliminaries
Pillar-based representation and sparse regional self-attention.
The pillar-based representation is an efficient representation introduced in PointPillars [23]. It divides 3D points into infinite-height voxels and computes pillar-wise features. These features can be treated as a pseudo-image in a bird’s eye view. On top of the pillar-based representation, SST [13] proposes a sparse regional self-attention (SRA) module to address a challenge in applying ViT to sparse LiDAR points. SST divides the pillars into non-overlapping windows with a fixed size and then applies self-attention to the non-empty pillars within the same window. The receptive field of each pillar is expanded through a window shift operation. T-MAE and many recent works [52, 48, 40] are built up on the basic block SRA.
Temporal batch.
A sweep of a point cloud at time $t$ is denoted by $\mathcal{P}^{t}=\{(p^{t}_{k})\}_{k=1}^{K}$, where $p_{k}$ is a 3D point $p_{k}\in\mathbb{R}^{3}$. Moreover, a sequence of point clouds is an ordered set of $\mathcal{P}^{t}$, which is denoted by $\mathcal{P}=\{\mathcal{P}^{1},\mathcal{P}^{2},\ldots,\mathcal{P}^{T}\}$. During pre-training and fine-tuning, two scans are needed feed the Siamese encoder. Due to the high sampling frequency of the LiDAR system, two consecutive frames usually contain redundant duplicate information, which is verified in Sec. 5. Conversely, when there is a substantial time gap between two frames, their overlap may be limited, rendering the historical information less useful. Therefore, we follow TARL [33] to introduce the concept of a temporal batch so that the interval of two frames is constrained to an appropriate range. Specifically, we sample a batch of consecutive frames $\mathcal{B}^{t}=\{\mathcal{P}^{t+1},\mathcal{P}^{t+2},\ldots,\mathcal{P}^{t+n}\}$ from a sequence $\mathcal{P}$. Then, two frames, $\mathcal{P}^{t+t_{1}}$ and $\mathcal{P}^{t+t_{2}}$ are sampled from the first and last one-third of the batch. More specifically, $0\leq t_{1}<\frac{n}{3}$ and $\frac{2n}{3}\leq t_{2}$.
3.2 Temporal Masked Autoencoder Pre-training
Inspired by the SiamMAE [17] used in video understanding, we develop Temporal Masked Auto-Encoders, namely T-MAE, for self-supervised learning on LiDAR points. As shown in Fig. 2, the core idea is to reconstruct the present frame based on the historical and partial observations of the current frame. In this way, the network is compelled to learn a powerful sparse representation as well as the capacity to effectively model motion. After the pre-training, the weights of the SiamWCA backbone are retained for the downstream tasks. In the subsequent subsections, we provide a detailed elaboration of these steps.
Single-frame baseline.
To perform the reconstruction pretext task, we follow the concept of the state-of-the-art SSL method GD-MAE [52] which operates on a single frame. The pipeline encompasses several key stages, including voxelization, masking, encoding, decoding, and reconstruction. More specifically, the initial step voxelizes a given point cloud by two linear layers and a voxel-wise average pooling. The voxel representation can be considered a pillar-based representation, owing to its infinite height. Next, masking is applied in a pillar-wise manner. Subsequently, the remaining visible pillars serve as input for the sparse pyramid transformer to generate high-level tokens. Next, the sparse tokens are transformed into a dense tensor by means of zero-padding. Subsequently, a series of dense convolutional layers are utilized to dilate the tokens of unmasked voxels to vacant regions.
$t$ ${t-3}\ \sim\ t$
Refer to caption Refer to caption
(a) A stationary vehicle.
Refer to caption Refer to caption (b) A moving vehicle.
Figure 3: Comparison between single and four frame concatenation. While simple frame concatenation generally improves point density and detection rates, it can introduce spurious points in non-static scene parts that may degrade the detection performance. Since we combine consecutive frames via learned cross-attention, our approach is less affected by this problem. The blue bounding boxes indicate the ground truth for the current frame.
Figure 4: Window-based sparse cross-attention (WCA) module. Given the input tokens from both the previous and current frames, a joint region grouping is conducted to obtain a window partition. A sparse regional cross-attention (SRCA) is performed individually in each window to integrate the historical information to the middle tokens for the current frame. In other words, the tokens from two frames but with the same colors are attending to each other. For simplicity, the information flow is only illustrated for the green tokens. The grouping and cross-attention are performed once more with the shifted window partition. The red dot (Refer to caption) indicates the ego-vehicle driving towards the right. The box with diagonal stripes (Refer to caption) represents an object moving towards the left. The observed scene remains mostly static except for a pillar at bottom-left transitions from occupied to vacant due to occlusion and another pillar which does the opposite. Best viewed in color and high-resolution.
SiamWCA.
Up to this point, a common approach has been to employ a single-frame reconstruction pretext task for LiDAR points. However, incorporating historical frames poses a non-trivial challenge. One straightforward approach is to concatenate two point clouds after aligning them with ego-poses, similar to the three-frame variant of SST [13]. However, as shown in Fig. 3, while concatenation helps to identify static objects effectively, it introduces challenges for detectors when dealing with moving objects. We provide quantitative results in Tab. 2. To address this problem, we suggest learning from past data in a feature space.
Given two frames, the previous frame is transformed to the coordinate system of the later frame by their ego-poses. Pose information is usually readily available in relevant datasets [39, 2, 29] or easily obtained by means of GPS/IMU, odometry approaches [8], structure from motion (SfM) algorithms [37, 38], or SLAM systems [10]. As shown in Fig. 2, the full observation of the previous frame is given to the network by passing it through voxelization and encoder. On the contrary, the later frame goes through voxelization, masking, and encoder. The encoders for the two frames share weights to ensure that the features are constrained in the same space.
By utilizing the Siamese encoder, tokens are obtained for both frames. These tokens serve as input to the WCA module, which facilitates the interaction between historical and current tokens and is outlined in Sec. 3.3. Once the current tokens are augmented with historical information, these sparse tokens are placed back into the x-y plane to form a dense tensor where vacant pillars are padded with zeros. Subsequently, a group of $3\times 3$ convolutional layers is utilized to process the tensor, which diffuses the features of visible pillars. Eventually, the obtained tokens of masked pillars are employed for point reconstruction. The Chamfer distance is computed as the loss function between the reconstructed points and the ground-truth points. Note that the number of points per pillar varies drastically. Thus, a fixed number of points $K^{GT}$ are randomly sampled as the target for reconstruction.
In conclusion, SiamWCA, containing a Siamese encoder and a WCA module, computes the pillar-based representation for both frames and integrates the historical one to the present one for reconstruction. Unlike the single-frame baseline where only the encoder is reused for downstream tasks, the weights of SiamWCA are fully retained for downstream tasks.
3.3 Window-based Sparse Cross-Attention Module
When provided with input tokens from two frames, there are several methods for interaction. One intuitive approach is to apply a standard self-attention layer to the concatenation of all tokens from both frames. This approach significantly increases GPU memory requirements. The vanilla Transformer block [43], which is combined with a cross-attention layer and a self-attention layer, can also be adopted. This combination, however, is demonstrated to be suboptimal as outlined in Sec. 4.3. In due course, a single cross-attention layer is evaluated and exhibits superior performance compared to others.
However, conventional global attention is not affordable in 3D space due to the expensive computational overhead imposed by the unavoidable high resolution. Therefore, the use of window-based attention is vital. The WCA module is illustrated in Fig. 4. The simulation involves an ego-vehicle moving to the right and another object moving in the opposite direction. The scene remains mostly static, but a pillar at bottom-left transitions from occupied to vacant due to occlusion.
Joint Regional Grouping
organizes the 3D spaces into non-overlapping windows. Hence, the following attention mechanism only performs with tokens from the same local region. The tokens from two frames are jointly considered so that the local windows across the two frames are corresponding. As shown in Fig. 4, all tokens within the same physical window share the same color, indicating that they are assigned to the same group for mutual attention.
Sparse Regional Cross-Attention
(SRCA) is essentially a cross-attention layer where the query comes from the current frame and the key-value comes from the previous frame. For a group of tokens $\mathcal{F}^{t_{1}},\mathcal{F}^{t_{2}}$ and their corresponding spatial coordinates $\mathcal{I}^{t_{1}},\mathcal{I}^{t_{2}}$, the query $\mathcal{Q}$, key $\mathcal{K}$, and value $\mathcal{V}$ are calculated as follows
$\mathcal{Q}$ $=\mathcal{F}^{t_{2}}+\mathbf{PE}(\mathcal{I}^{t_{2}})$ (1)
$\mathcal{K}$ $=\mathcal{F}^{t_{1}}+\mathbf{PE}(\mathcal{I}^{t_{1}})$ (2)
$\mathcal{V}$ $=\mathcal{F}^{t_{1}}$ (3)
where $\mathbf{PE}(\cdot)$ represents the absolute positional encoding function used in [4]. SRCA follows conventional cross-attention and is defined as follows
^$\hat{\mathcal{F}}^{t_{2}}$ $=\mathbf{MCA}(\mathcal{Q},\mathcal{K},\mathcal{V})$ (4)
~$\widetilde{\mathcal{F}}^{t_{2}}$ $=\mathbf{LN}\left(\mathbf{MLP}\big{(}\mathbf{LN}(\hat{\mathcal{F}}^{t_{2}})\big{)}+\hat{\mathcal{F}}^{t_{2}}\right)$ (5)
where $\mathbf{MCA}(\cdot,\cdot,\cdot)$ indicates the Multi-head Cross-Attention, and $\mathbf{LN}(\cdot)$ stands for Layer Normalization. While the core concept is similar to cross-attention, this implementation provides a more efficient solution when working with sparse pillars.
Repeated Operations with Region Shift.
The joint regional grouping and SRCA are performed once more after the regions are shifted. We follow the implementation of SST [13] to achieve region shift. Formally, suppose that the window size is $(l_{x},l_{y},l_{z})$, then the original windows are moved by $(l_{x}/2,l_{y}/2,l_{z}/2)$. Meanwhile, the tokens are re-grouped by the joint regional grouping, as illustrated in “Shifted Window Partition” of Fig. 4. Note that, while performing the SRCA #2, the tokens ~$\widetilde{\mathcal{F}}^{t_{2}}$ of the current frame are updated by SRCA #1 while the tokens $\mathcal{F}^{t_{1}}$ remain unchanged.
4 Experiments
4.1 Dataset and Implementation
Waymo Open dataset [39] is the currently the largest public dataset with LiDAR points for autonomous driving. It comprises 798 sequences for training and 202 sequences for validation. Each sequence contains roughly 200 frames with pose information by which points can be transformed into the world coordinate system. For 3D detection, an evaluation protocol is provided for calculating the average precision (AP) and the average precision weighted by heading (APH). Moreover, the evaluation includes two difficulty levels wherein bounding boxes containing over 5 points are regarded as Level 1 while Level 2 indicates all bounding boxes. We adopt mean AP and mean APH at Level 2 as main evaluation metrics.
Implementation Details. We implement our approach based on the codebase of OpenPCDet [^1] https://github.com/open-mmlab/OpenPCDet. We keep the configuration of the sparse pyramid transformer consistent with its official implementation of GD-MAE [52]. The window size of the WCA modules is set to $(8,8,1)$. The detection ranges for the Waymo dataset are restricted between $(-74.88m,-74.88m,-2m)$ and $(74.88m,74.88m,4m)$. The pillar size is $(0.32m,0.32m,6m)$. The size of the temporal batch is set to $n=6$ for both pre-training and finetuning. Three data augmentation techniques, i.e. random flipping, scaling, and rotation, are applied to both frames during pre-training and finetuning. Note that, the variables of augmentations applied to two frames keep the identical. We use the AdamW optimizer [28] with the one-cycle policy to train the model. The max learning rate is set to $3e^{-3}$. All experiments are performed using 8 NVIDIA Tesla V100.
Pre-training. The model is pre-trained for 12 epochs on the full training split of Waymo dataset, which contains 158,081 frames. The batch size is set to 4. Following GD-MAE [52], the masking ratio is set to 0.75, the number of predicted points per pillar is set as 16 and the number of target reconstruction points is set as $K^{GT}=64$.
Finetuning and inference. The model is finetuned for 30 epochs on different subsets of the Waymo training split. The number of epochs is held constant at 30, with a batch size of 2, while the data quantity varies. In addition to the data augmentations mentioned in the implementation details of Sec 4.1, following MV-JAR [48] and GD-MAE [52], a copy-n-paste augmentation [51] is also employed to slightly address the issue of class imbalance. During inference, each frame is selected as a current frame for a complete evaluation. The previous third frame, based on our ablation study, is selected as the previous frame.
4.2 Results on Waymo
In this section, the main results of the Waymo dataset are presented. We assess our approach after fine-tuning SiamWCA with varying amounts of labeled data. We explore two methods to obtain a specific percentage of the training set, but all models are evaluated on the same validation set.
$\triangleright$ Uniform sampling is commonly used to obtain a subset of LiDAR sequence [53, 52, 25]. Specifically, frames from all sequences are concatenated, from which the target frames are sampled with a fixed interval.
Data Amount Initialization Overall Vehicle Pedestrian Cyclist
L2 mAP L2 mAPH L2 mAP L2 mAPH L2 mAP L2 mAPH L2 mAP L2 mAPH
5% Random 43.68 40.29 54.05 53.50 53.45 44.76 23.54 22.61
PointContrast [47] 45.32 41.30 52.12 51.61 53.68 43.22 30.16 29.09
ProposalContrast [53] 46.62 42.58 52.67 52.19 54.31 43.82 32.87 31.72
MV-JAR [48] 50.52 46.68 56.47 56.01 57.65 47.69 37.44 36.33
GD-MAE [$*{}^{*}$ 48.23 44.56 56.34 55.76 55.62 46.22 32.72 31.69
T-MAE (Ours) 50.20${}^{+6.52}$ 48.07${}^{+7.78}$ 56.71 56.18 60.19 55.71 33.69 32.33
10% Random 56.05 53.13 59.78 59.27 60.08 53.04 48.28 47.08
PointContrast [47] 53.69 49.94 54.76 54.30 59.75 50.12 46.57 45.39
ProposalContrast [53] 53.89 50.13 55.18 54.71 60.01 50.39 46.48 45.28
MV-JAR [48] 57.44 54.06 58.43 58.00 63.28 54.66 50.63 49.52
GD-MAE [$*{}^{*}$ 57.67 54.31 59.72 59.19 60.43 52.21 52.85 51.52
T-MAE (Ours) 59.14 ${}^{+3.09}$ 57.05 ${}^{+3.92}$ 60.29 59.79 63.61 59.24 53.52 52.13
20% Random 60.21 57.61 61.58 61.08 64.63 58.41 54.42 53.33
PointContrast [47] 59.35 55.78 58.64 58.18 64.39 55.43 55.02 53.73
ProposalContrast [53] 59.52 55.91 58.69 58.22 64.53 55.45 55.36 54.07
MV-JAR [48] 62.28 59.15 61.88 61.45 66.98 59.02 57.98 57.00
GD-MAE [$*{}^{*}$ 62.32 59.09 62.27 61.79 66.12 58.06 58.57 57.42
T-MAE (Ours) 63.48 ${}^{+3.27}$ 61.72 ${}^{+4.11}$ 62.93 62.47 68.26 64.53 59.24 58.16
20% Uniform Sampling Random 69.03 66.48 66.61 66.15 72.65 66.76 67.83 66.54
ProposalContrast [53] 66.67 64.20 65.22 64.80 66.40 60.49 68.48 67.38
BEV-MAE [26] 66.70 64.25 64.72 64.22 66.20 60.59 69.11 67.93
Occupancy-MAE [31] 65.86 63.23 64.05 63.53 65.78 59.62 67.76 66.53
MAELi-MAE [22] 65.60 63.00 64.22 63.70 65.93 59.79 66.66 65.52
GD-MAE [${}^{**}$ 70.24 67.14 67.67 67.22 73.18 65.50 69.87 68.71
AD-PT [55] 67.17 64.65 65.33 64.83 67.16 61.20 69.39 68.25
T-MAE (Ours) 70.61 ${}^{+1.58}$ 68.80 ${}^{+2.32}$ 66.85 66.39 74.85 71.08 70.14 68.94
100% Random 71.30 69.13 69.05 68.62 73.77 68.80 71.09 69.97
GCC-3D [25] 65.29 62.79 63.97 63.47 64.23 58.47 67.88 66.44
BEV-MAE [26] 66.92 64.45 64.78 64.29 66.25 60.53 69.73 68.52
PointContrast [47] 68.06 64.84 64.24 63.82 71.92 63.81 68.03 66.89
ProposalContrast [53] 68.17 65.01 64.42 64.00 71.94 63.94 68.16 67.10
MV-JAR [48] 69.16 66.20 65.52 65.12 72.77 65.28 69.19 68.20
GD-MAE [${}^{**}$ 70.62 67.64 68.72 68.29 72.84 65.47 70.30 69.16
T-MAE (Ours) 71.96 ${}^{+0.65}$ 70.22 ${}^{+1.09}$ 69.07 68.63 74.74 71.03 72.06 70.99
Table 1: Performance comparison with SSL methods on the Waymo validation set [39]. Random initialization denotes training frame scratch. 20% Uniform Sampling indicates sampling frames per five frames. $*{}^{*}$ indicates reproduced by us. ${}^{**}$ indicates from [52]. Results for other methods are taken from MV-JAR [48] or the survey [14]. Best results are highlighted as first, second, and third. Differences between T-MAE pre-training and random initialization are highlighted in red. T-MAE achieves significant improvements compared to random initialization and SOTA results compared to other SSL methods.
$\triangleright$ Data-efficient benchmark is proposed in MV-JAR [48] to solve the data diversity issue existing in the strategy of uniform sampling. More specifically, while trained on uniformly sampled frames, the models always converge to a similar performance if they are trained for adequate iterations no matter how many percentages of labeled data are used. As a result, fewer data do not lead to a shorter finetuning time. Therefore, to evaluate pre-trained representation more efficiently, we conduct experiments on the data-efficient benchmark, as presented in Tab. 1. Note that, for each data amount level, three non-overlap splits are provided MV-JAR [48]. The results for models finetuned by the second or the third split are presented in the supplementary material, as well as the average performance.
$\triangleright$ T-MAE achieves new SOTA at labelled data level of 10%, 20%, 20% (Uniform sampling) and 100%. Moreover, T-MAE achieves superior performance for the majority of class-specific metrics. To reduce performance variance on 5% and 10% splits, we also report in the supplementary material the results of T-MAE by finetuning it on the other two splits provided by [52]. Notably, thanks to the temporal modeling ability, T-MAE significantly outperforms other methods in terms of L2 mAPH for pedestrians. For instance, with 5% labeled data, T-MAE achieve better mAPH (i.e. 55.71) than any other method using 10% labeled data. The comparison with other SSL methods shows the efficacy of T-MAE. Nevertheless, MV-JAR [48] performs notably better for cyclists at 5% level and surpasses T-MAE in terms of L2 overall mAP.
$\triangleright$ The impact of the T-MAE pre-training becomes more pronounced when the labeled data is scarce. SiamWCA is a powerful backbone as, when finetuned with 100% of labeled data, it even performs better than other models that are initialized by different pre-training strategy (i.e. 69.13 v.s. 67.64). However, its performance drops significantly when finetuning data is limited (i.e. 40.29 v.s. 46.68 at 5% level). T-MAE consistently enhances model performance compared to random initialization, with more significant improvements observed when labeled data is limited.
4.3 Ablation study
In this section, we perform a series of ablation studies to assess the necessity of various components and justify the selection of hyperparameters. To cost-effectively experiment, the split 0 with 5% data of the data-efficient benchmark is used to finetune the model by default.
Is a delicate fusion module necessary? To compare with a model taking two frames as input, we modified the input of GD-MAE [52] by concatenating all points of two consecutive frames. We pre-trained, finetuned, and evaluated the modified GD-MAE with the same setting as T-MAE. As depicted in Table 2, directly merging two frames improves the metrics in terms of vehicles and pedestrians, which probably results from the density of target objects being doubled, as shown in Fig. 3. However, it has a negative impact on cyclists, resulting in a decrease in overall performance. The drop for cyclists might be attributed to the drift of estimated bounding boxes caused by the fast velocity of bicycles and their relatively small dimensions. In contrast, T-MAE significantly improves the performance thanks to the delicate pre-training strategy and fusion mechanism.
Method Frame Input L2 Overall Vehicle Pedestrian Cyclist
Previous Current mAP mAPH AP APH AP APH AP APH
GD-MAE - 0 48.23 44.56 56.34 55.76 55.62 46.22 32.72 31.69
GD-MAE - {-1, 0} 47.35 43.69 57.04 56.48 57.76 48.48 27.26 26.12
T-MAE (Ours) -1 0 49.45 46.78 56.56 56.02 57.96 51.94 33.82 32.37
Table 2: Two frames comparison. Two consecutive frames are merged and input into the GD-MAE [52] as an enhanced baseline.
Model Encoder Fusion L2 Overall Vehicle Pedestrian Cyclist
mAP mAPH AP APH AP APH AP APH
(a) Asymmetric WCA 47.26 44.78 54.25 53.75 56.66 50.98 30.88 29.60
(b) SimSiam WCA 45.58 42.05 53.37 52.84 53.80 44.78 29.58 28.53
(c) Siamese WCA$+$WSA 47.01 45.11 55.81 55.33 58.95 54.69 26.26 25.31
(d) Siamese ST WCA$+$WSA 46.29 43.64 54.13 53.64 55.73 49.24 29.02 28.04
(e) Ours Siamese WCA 49.45 46.78 56.56 56.02 57.96 51.94 33.82 32.37
Table 3: Ablation study on model architecture. The proposed SiamWCA demonstrates superior performance in terms of overall mAP and mAPH.
Encoder design and fusion mechanism. The proposed architecture consists of a Siamese encoder and a WCA module, which is denoted as (e) in this ablation study. In addition, several variants are compared: (a) Asymmetric encoder: The encoder for the previous frame is replaced with a more lightweight encoder, resulting in a reduction of the number of channels by half. Consequently, the two encoders no longer share weights. (b) We employ the SimSiam [7] approach, where one of the two encoders is disconnected from the computational graph, preventing it from being updated by gradient propagation. The encoder for the current frame is updated by gradients and then shares its weights with the detached encoder. (c) The design of the Siamese encoder stays unaltered, while a typical cross-self attention module is adopted in a window-based manner. These three variants are depicted in Fig. 5. (d) Instead of using GD-MAE [52] as our single-frame backbone, the standard (ST) MAE [18] is adopted to 3D space and serves as the Siamese encoder. The superiority of the proposed SiamWCA is shown in Tab. 3. Variant (c) outperforms SiamWCA in terms of pedestrians but at the expense of cyclists’ performance.
(a) Asym. Enc. (b) SimSiam Enc. (c) WCA + WSA
Figure 5: Three architecture variants of our SiamWCA backbone. (a) Asymmetric encoders: the encoder for the previous frame is replaced by a more lightweight encoder. (b) SimSiam-style [7] encoder: one encoder receives no gradient updates. (c) An additional windows-based self-attention is attached as the classic Transformer [43].
Previous frame L2 Overall Vehicle Pedestrian Cyclist
mAP mAPH AP APH AP APH AP APH
$[-5,-1]$ 49.86 47.61 56.70 56.16 59.47 54.65 33.41 32.03
$-1$ 49.45 46.78 56.56 56.02 57.96 51.94 33.82 32.37
$-3$ 50.20 48.07 56.71 56.18 60.19 55.71 33.69 32.33
$-5$ 49.87 47.69 56.73 56.20 59.91 55.26 32.98 31.61
Table 4: Temporal gap for inference. Different temporal gaps were tested for inference. $[-5,-1]$ indicates a random frame is sampled from the last five frames. Overall, a three-frame temporal gap worked best in our experiments.
Temporal interval for inference. This ablation study explores the impact of different temporal gaps during inference on model performance. We establish a baseline with a random interval between two frames. Specifically, for each frame, the previous frame is selected randomly from the preceding five frames. Furthermore, the temporal gap is fixed to $-1$, $-3$, and $-5$, respectively. If the number of available frames is less than the predefined threshold, the earliest available frame is selected. Tab. 4 illustrates that optimal performance is achieved with a fixed temporal gap of three. Specifically, the model excels in recognizing vehicles, pedestrians, and cyclists when the temporal gap is set to $-5$, $-3$, and $-1$, respectively. This observation highlights the varied impact of different temporal gaps on distinct object categories. Notably, while the performance remains relatively stable for vehicles and cyclists, there is a pronounced improvement in pedestrian recognition as the temporal gap transitions from $-1$ to $-3$. This highlights the considerable advantage of including more distant historical frames to achieve accurate pedestrian detection as well as his/her heading.
5 Conclusion
In this paper, we introduced the Temporal Mask AutoEncoders (T-MAE), a novel self-supervised paradigm for LiDAR point cloud pre-training. Building upon the single-frame MAE baseline, we incorporated historical frames into the representation using SiamWCA, with the proposed WCA module playing a pivotal role. This pre-training enabled the model to acquire robust representations and the ability to capture motion even with very limited labeled data. By considering temporally distant frames instead of consecutive ones, we achieved additional performance improvements. Our experiments on the Waymo Open dataset demonstrate the effectiveness of our approach by showing mostly improvements over state-of-the-art methods.
6 Acknowledgements
This work was financially supported by TomTom, the University of Amsterdam and the allowance of Top consortia for Knowledge and Innovation (TKIs) from the Netherlands Ministry of Economic Affairs and Climate Policy. Fatemeh Karimi Nejadasl was financed by the University of Amsterdam Data Science Centre.
Data Initialization Overall Vehicle Pedestrian Cyclist
amount L2 mAP L2 mAPH L2 mAP L2 mAPH L2 mAP L2 mAPH L2 mAP L2 mAPH
5% Random 49.59 46.15 54.42 53.87 52.59 44.17 41.76 40.41
PointContrast [47] 48.97 44.91 52.35 51.85 52.49 41.95 42.07 40.91
ProposalContrast [53] 49.87 45.83 52.79 52.31 53.30 43.00 43.51 42.18
MV-JAR [48] 52.73 48.99 56.66 56.21 57.52 47.61 44.02 43.15
T-MAE (Ours) 52.91 ${}^{+3.32}$ 50.98 ${}^{+4.83}$ 55.60 55.05 57.98 54.18 45.14 43.70
10% Random 57.44 54.48 59.63 59.10 60.38 53.25 52.31 51.09
PointContrast [47] 55.22 51.31 55.62 55.15 59.25 49.17 50.81 49.60
ProposalContrast [53] 55.59 51.67 55.57 55.12 60.02 49.98 51.18 49.90
MV-JAR [48] 58.61 55.12 58.92 58.49 63.44 54.40 53.48 52.47
T-MAE (Ours) 59.18 ${}^{+1.74}$ 57.29 ${}^{+2.81}$ 59.66 59.15 63.90 59.87 53.97 52.84
Table 5: Performance comparison on the Waymo validation set [39]. Results for other methods are taken from MV-JAR [48]. All methods are finetuned with Subset 1.
Data Initialization Overall Vehicle Pedestrian Cyclist
amount L2 mAP L2 mAPH L2 mAP L2 mAPH L2 mAP L2 mAPH L2 mAP L2 mAPH
5% Random 43.33 40.10 55.39 54.81 55.75 47.69 18.86 17.79
PointContrast [47] 44.48 40.55 51.87 51.37 55.36 45.03 26.22 25.24
ProposalContrast [53] 45.21 41.45 52.29 51.82 56.23 46.28 27.10 26.24
MV-JAR [48] 47.93 44.50 56.22 55.78 58.80 49.77 28.75 27.95
T-MAE (Ours) 46.89 ${}^{+3.56}$ 45.04 ${}^{+4.95}$ 56.32 55.81 60.90 56.65 23.45 22.67
10% Random 56.81 53.97 59.49 58.98 62.44 55.54 48.50 47.40
PointContrast [47] 54.80 51.02 55.41 54.95 60.56 50.86 48.44 47.24
ProposalContrast [53] 54.77 51.09 55.64 55.20 60.54 51.16 48.14 46.92
MV-JAR [48] 58.29 54.99 59.17 58.74 64.58 56.02 51.12 50.20
T-MAE (Ours) 58.15 ${}^{+1.34}$ 56.26 ${}^{+2.29}$ 60.08 59.56 66.02 61.82 48.34 47.41
Table 6: Performance comparison on the Waymo validation set [39]. Results for other methods are taken from MV-JAR [48]. All methods are finetuned with Subset 2.
Data Initialization Overall Vehicle Pedestrian Cyclist
amount L2 mAP L2 mAPH L2 mAP L2 mAPH L2 mAP L2 mAPH L2 mAP L2 mAPH
5% Random 45.53 42.18 54.62 54.06 53.93 45.54 28.05 26.94
PointContrast [47] 46.26 42.25 52.11 51.61 53.84 43.40 32.82 31.74
ProposalContrast [53] 47.23 43.28 52.58 52.10 54.61 44.37 34.50 33.38
MV-JAR [48] 50.39 46.72 56.45 56.00 57.99 48.36 36.74 35.81
T-MAE (Ours) 50.00 ${}^{+4.46}$ 48.03 ${}^{+5.85}$ 56.21 55.68 59.69 55.51 34.09 32.90
10% Random 56.77 53.86 59.63 59.12 60.97 53.94 49.70 48.52
PointContrast [47] 54.57 50.75 55.26 54.80 59.85 50.05 48.61 47.41
ProposalContrast [53] 54.75 50.96 55.47 55.01 60.19 50.51 48.60 47.37
MV-JAR [48] 58.12 54.72 58.84 58.41 63.77 55.03 51.74 50.73
T-MAE (Ours) 58.82 ${}^{+2.06}$ 56.87 ${}^{+3.01}$ 60.01 59.50 64.51 60.31 51.94 50.79
Table 7: Average performance on the Waymo validation set [39], averaged across the models finetuned with Subset 0--2. Results for other methods are taken from MV-JAR [48]. T-MAE consistently enhances model performance with a notable margin compared to random initialization. Furthermore, T-MAE attains SOTA average performance across all metrics when finetuned with 10% data.
(a) Input (b) Attention scores for prior scan
Figure 6: Visualization of attention scores. (a) The input consists of two point clouds: the entire previous scan (red points) and the current scan (green points) that contains only points within the ground-truth bounding boxes. Note that, the blue bounding boxes in (a) serve solely for visualization purposes and do not function as input. (b) The pillar-wise attention scores are visualized. The attention scores are derived from the WCA module and mapped to a colormap ranging from black to white. The primary attention is placed on the target objects from the previous scan. This implies that the WCA is able to locate corresponding objects. Notably, WCA is capable of accurately trace to the source object, which is manually indicated by a skyblue box, even in cases where the vehicle is moving.
Method Data amount Epochs Number of input scans (×$\times 10^{3}$ per epoch) Total iterations (×$\times 10^{5}$) Number of GT used (×$\times 10^{3}$ per epoch) L2 mAPH (Overall) L2 mAPH (Pedestrian)
MV-JAR [48] 5% 72 7.9 5.70 7.9 46.68 47.69
10% 60 15.8 9.50 15.8 54.06 54.66
20% 48 31.6 15.19 31.6 59.15 59.02
Random/T-MAE 5% 30 7.9 2.37 4.0 40.29/48.07 44.76/55.71
10% 30 15.8 4.75 7.9 53.13/57.05 53.04/59.24
20% 30 31.6 9.50 15.8 57.61/61.72 58.41/64.53
Table 8: Quantitative details on comparison of finetuning iterations. Our scheme requires much less iterations for finetuning compared to MV-JAR [48].
Appendix A Additional Results on Waymo Subsets
The data-efficient benchmark, proposed by MV-JAR [48], contains three splits, each of which contains four tiers of data, namely 5%, 10%, 20% and 50%. In the context of data organization, each split corresponds to a subdivision where a lower level of data is a subset of a higher level. There is no overlap observed between different splits for the data corresponding to the 5% and 10% subsets. In the main paper, the majority of experiments were performed by finetuning the model with the split 0. To alleviate the variance caused by being finetuned with a small amount of data, we further present the performance obtained when the model is finetuned using the remaining two splits, as shown in Tab. 5 and Tab. 6. Furthermore, the average results are presented in Tab. 7. In conclusion, it can be observed that the model performance exhibits variation when finetuned with different subsets. However, T-MAE consistently demonstrates a noteworthy improvement in model performance compared to random initialization. Additionally, T-MAE attains SOTA average performance across all metrics when finetuned using 10% data.
Appendix B Visualization of cross-attention scores
To understand the knowledge acquired by the SiamWCA during pre-training, we employ attention visualization to identify critical regions of the previous scan. In particular, the entire previous scan and the target objects of current scan are input into the SiamWCA. More precisely, the input to the model for the current scan solely consists of points contained within the 3D ground-truth bounding boxes. The purpose of removing other points is to identify the specific areas of emphasis within the WCA module when the queries are solely target objects, e.g. vehicles and pedestrians. As illustrated in Fig. 6 (b), the attention is mainly attached to target objects, indicating that the WCA successfully detects and localizes these entities in the previous frame. Furthermore, the presence of vehicular action is perceptible even if the WCA module is only trained with unlabeled data, indicating the effectiveness of our T-MAE pre-training strategy.
Appendix C Comparison of training iterations.
MV-JAR [48] applies varying numbers of epochs during finetuning, depending on the data amount. Rather than varying numbers, we employ a predetermined number of epochs, leading to significantly less iterations for finetuning without performance drop. Detailed information is presented in Tab. 8. The total number of iterations is calculated by multiplying the number of input scans by the number of epochs. As depicted in Fig. 1 and Tab. 8, T-MAE achieve better performance than MV-JAR [48] while requiring a smaller number of finetuning iterations. In addition, our scheme requires just one GT annotation every two input frames. As a result, the total number of iterations would be even smaller if it is calculated by multiplying the number of GT by the number of epochs.